Here's the question in my math book that I'm having problems with.
46. A guy wire to the top of a pole is 35 feet long. It reaches the ground 18 feet from the base of the pole. How tall is the pole?
I know that you have to use the theorem of pythagoras to figure it out, but i'm not sure which number goes where. Any explanation would be great.

Answer by checkley71(8403)

A^2+B^2=C^2
LET A=18 FEET
LET B=HEIGHT OF POLE
LET C=35 FEET GUIDE WIRE (HYPOTENUSE OF A RIGHT TRIANGLE)
18^2+B^2=35^2
144+B^2=1225
B^2=1225-144
B^2=1081
B=32.88 FEET FOR THE POLE